OpenStudy (anonymous):

Is the slope of this fraction 1/7 or just 7??? I just wanna be certain. Ha ha, don't need another 0, you know!!! ;D

OpenStudy (anonymous):

it would b 1/7 bc slope is rise over run so it should be up 1 over 7 if you are doing it correctly ;)

jimthompson5910 (jim_thompson5910):

would have to see the full problem to verify your answer

OpenStudy (anonymous):

Thank you, you smart person @britt0217 , and @jim_thompson5910 the full problem is "Find the slope of the line. The relations are (-2,2), (5,3)."

jimthompson5910 (jim_thompson5910):

Use the slope formula m = (y2-y1)/(x2-x1) m = (3-2)/(5-(-2)) m = (3-2)/(5+2) m = 1/7 So the slope of the line that passes through the two points is 1/7

jimthompson5910 (jim_thompson5910):

So your first answer was correct

OpenStudy (anonymous):

your right girl :)

OpenStudy (anonymous):

YES!! THANKS BROS!! @britt0217 and @jim_thompson5910 you guys are amazing!!! :D

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

your welcome :)